# Stepwise Alignment for Constrained Language Model Policy Optimization

### v1

@@ Abstract
Safety and trustworthiness are indispensable requirements for applying AI systems based on large language models (LLMs) in real-world applications. This paper formulates a human value alignment as a language model policy optimization problem to maximize reward under a safety constraint and then proposes an algorithm called Stepwise Alignment for Constrained Policy Optimization (SACPO). A key idea behind SACPO, supported by theory, is that the optimal policy incorporating both reward and safety can be directly obtained from a reward-aligned policy. Based on this key idea, SACPO aligns the LLMs with each metric step-wise while leveraging simple yet powerful alignment algorithms such as direct preference optimization (DPO). SACPO provides many benefits such as simplicity, stability, computational efficiency, and flexibility regarding algorithms and dataset selection. Under mild assumption, our theoretical analysis provides the upper bounds regarding near-optimality and safety constraint violation. Our experimental results show that SACPO can fine-tune Alpaca-7B better than the state-of-the-art method in terms of both helpfulness and harmlessness.
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities in diverse real-world applications [9] such as translation [45], content creation [44], coding [10, 16], summarization [37], robotics [35], among others. As the utilization of LLMs in artificial intelligence (AI) systems permeates our daily lives, the importance of responsible and ethical use grows; safety issues have been highlighted [17, 25, 27]. Consequently, as AI continues to evolve and become more integrated into society, it is crucial that we actively research and develop solutions to ensure that the benefits of AI are realized while minimizing any negative societal impacts.
To address this challenge, alignment has been used to embed human values and goals into LLMs to enhance their utility and safety. Notably, alignment premised on human feedback has emerged as a key mechanism in making LLMs more helpful and harmless, as represented by reinforcement learning from human feedback (RLHF, [11, 29]). Standard RLHF workflows fit a reward model to a human preference dataset and then optimize a language model (LM) policy to maximize the expected reward without diverging from the original policy. However, RLHF measures the quality of outputs in terms of a single metric (i.e., reward); thus, an achieved safety level is not usually high, and a model that refuses to answer renders the response quite unhelpful even if it is considered harmless [12].
@@
Safe RLHF [12] is a pioneering approach for introducing the (constrained) safe RL paradigm into the alignment of LLMs. As with standard RLHF, the safe RLHF pipeline trains reward and safety models from the human preference dataset and then employs RL to optimize an LM policy. This approach facilitates the acquisition of LLMs that strike a well-balanced compromise between reward (i.e., helpfulness) and safety (i.e., harmlessness). However, the Safe RLHF pipeline is inherently more complex than supervised learning or standard RLHF, as it necessitates fitting reward and safety models to preference data and then learning a policy using PPO Lagrangian [32] that simultaneously optimizes an additional parameter (i.e., Lagrangian multiplier) to balance helpfulness and harmlessness.
Our contributions. We propose an algorithm called Stepwise Alignment for Constrained Policy Optimization (SACPO) for human value alignment of LLMs while incorporating decoupled reward and safety metrics. As shown in Figure 1, SACPO is a stepwise approach that sequentially aligns an LLM with one metric (e.g., reward) and subsequently with another (e.g., safety). Our theoretical analysis allows us to employ simple RL-free alignment algorithms such as direct preference optimization (DPO, [31]) or Kahneman-Tversky optimization (KTO, [15]) for each alignment without necessitating explicit reward or safety modeling. In a theoretically justified way, SACPO allows us to use different alignment algorithms or parameters for each alignment, thereby enhancing the flexibility of the format or volume of the datasets. To augment the practicality of SACPO, we further propose an efficient approach using model merging to balance the trade-off between reward and safety performance. We provide theoretical results regarding the optimality and safety of the LM policy obtained by SACPO under mild assumptions. Finally, our experimental results show that SACPO can fine-tune Alpaca-7B better than the Safe RLHF in terms of both helpfulness and harmlessness.
@@ 2 Preliminaries
Given a pre-trained LLM, the alignment is traditionally conducted in two stages [5, 29, 48]. In the first stage, called supervised finetuning (SFT), a pre-trained LLM is fine-tuned with a cross-entropy loss over the high-quality human completion and then obtains a model $\pi_{\text{SFT}}$. This stage enables the model to predict the next token more properly on data more relevant for downstream tasks (e.g., dialogue, summarization). The second stage (i.e., learning from human feedback) aims better to align LLMs to human desiderata [11]. Because this paper focuses on the second stage, we will review the existing representative approaches and algorithms for the second stage. Specifically, we will briefly review RLHF [29] and then explain subsequent RL-free approaches such as DPO [31] or KTO [15].
@@ 2.1 Reinforcement Learning from Human Feedback (RLHF)
The RLHF pipeline consists of the following two phases: 1) reward modeling and 2) RL finetuning phases. With a prompt $x\in\mathcal{X}$, an LLM is regarded as a (stochastic) policy to return an output $y\in\mathcal{Y}$, where $\mathcal{X}$ and $\mathcal{Y}$ are respectively the finite spaces of prompts and outputs. Here, we assume access to a dataset of preferences $\mathcal{D}\coloneqq\{(x^{(i)},y_{w}^{(i)},y_{l}^{(i)})\}_{i=1}^{N}$, where $y_{w}$ and $y_{l}$ denote preferred and dispreferred outputs (i.e., $y_{w}\succ y_{l}$) for a prompt $x$. Paired outputs $(y_{w},y_{l})$ are typically sampled from $\pi_{\text{SFT}}$.
Reward modeling. In the first stage, the preference dataset $\mathcal{D}$ is assumed to be generated by a latent (unknown) reward model $r^{\star}$, and a typical choice is the Bradley-Terry (BT) model [8] stipulating that the human preference distribution $p^{\star}$ is written as
@@
RL finetuning. In the second stage, we aim to maximize the reward while leveraging a KL divergence penalty to restrict how far the language model can drift from the base reference policy $\pi_{\mathrm{ref}}$, namely the SFT model $\pi_{\text{SFT}}$. Let $\pi_{\theta}$ denote the model we are optimizing and we solve the following policy optimization problem to maximize
@@
where $\rho$ is a distribution of prompts used in the RL phase, and $\mathbb{E}_{\rho,\pi}[\cdot]$ is an abbreviated notation for $\mathbb{E}_{x\sim\rho,y\sim\pi(\cdot\mid x)}[\cdot]$ for any policy $\pi\in\Pi$. Note that $\beta>0$ is a hyperparameter for the KL penalty. Since this objective is not differentiable, we leverage RL algorithms such as PPO [34] as in Ouyang et al. [29] or REINFORCE [41] as in Ahmadian et al. [1] to optimize it.
@@ 2.2 Direct Learning from Human Feedback without RL
RLHF (especially when based on PPO) is slow and unstable in practice; thus, many algorithms have been recently proposed to overcome this issue. A common idea is to analytically derive the optimal policy of (3) and parameterize it using the reward function as follows:
@@
Here, for any function $f:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ and policy $\widetilde{\pi}\in\Pi$, $Z_{f}(x;\widetilde{\pi})$ is a normalization term or constant defined as $Z_{f}(x;\widetilde{\pi})\coloneqq\sum_{y}\widetilde{\pi}(y\mid x)\exp\left(\frac{1}{\beta}f(x,y)\right)$. For the proof, see Appendix A. Note that, to derive the optimal policy (4), we do not assume a specific reward structure such as the BT model. Thus, the overall structure of the optimal policy results from the problem setting (10) characterized by the reverse KL, which is common to the representative algorithms listed below.
@@
Since $Z_{r}(x;\pi_{\mathrm{ref}})$ does neither depend on $y$ nor $\pi$, by simply plugging the reward (5) into the loss function associated with the BT model (2), the resulting objective of DPO is given by
@@
As a generalized extension of DPO, Azar et al. [4] proposed $\Psi$PO characterized by a more general objective function, encompassing those of RLHF and DPO, which is exclusively expressed in terms of pairwise preferences. IPO is proposed as a specific case of $\Psi$PO to bound the reward function for avoiding overfitting where the loss function is defined as:
$\mathcal{L}_{\text{IPO}}(\pi_{\theta},\pi_{\text{ref}},\beta)=\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\,\left(\log\left(\frac{\pi_{\theta}(y_{w}\mid x)\pi_{\mathrm{ref}}(y_{l}\mid x)}{\pi_{\theta}(y_{l}\mid x)\pi_{\mathrm{ref}}(y_{w}\mid x)}\right)-\frac{1}{2\beta}\right)^{2}\,\right].$ (7)
KTO. The algorithms discussed above need a preference dataset, which is costly for humans to collect. Ethayarajh et al. [15] propose Kahneman-Tversky optimization (KTO) that needs only a binary signal of whether the output $y$ is desirable (i.e., $y_{+}$) or undesirable (i.e., $y_{-}$) for a given prompt $x$. With an unpaired dataset $\widetilde{\mathcal{D}}\coloneqq\{(x^{(i)},y^{(i)})\}_{i=1}^{N}$, the loss function for KTO is calculated as
@@
where $v_{\text{KTO}}$ is called a value function to maps a latent reward $r_{\text{KTO}}(x,y)\coloneqq\beta\log{\frac{\pi_{\theta}(y\mid x)}{\pi_{\text{ref}}(y\mid x)}}$, relative to some reference point $\nu\coloneqq\beta\,\mathbb{D}_{\mathrm{KL}}[\,\pi_{\theta}(y\mid x)\,\|\,\pi_{\text{ref}}(y\mid x)\,]$, to its perceived value:
@@ 2.3 Safe and Multi-objective Learning from Human Feedback
Although all the algorithms discussed above incorporate only a singular reward function, several algorithms with constraints or multiple objectives have been proposed in recent years [47, 12, 46, 28].
Safe RLHF. To improve the helpfulness and harmlessness of LLMs, Dai et al. [12] additionally introduce a safety function $g^{\star}$ and then formulate an LLM alignment problem as a policy optimization task of maximizing the reward $r^{\star}$ while satisfying a safety constraint. They propose Safe RLHF that first trains reward and safety models (i.e., $r_{\psi}$ and $g_{\psi}$) using two datasets containing reward and safety information, and then optimizes the policy by solving the following problem using a popular safe RL algorithm called PPO-Lagrangian [32]:
@@
Safe RLHF requires us to fit reward and safety models and then concurrently optimize the LM policy and Lagrangian multiplier to balance helpfulness and harmlessness. Hence, Safe RLHF is a more complex and unstable procedure, even when compared to standard RLHF.
Multi-objective and constrained DPO. Zhou et al. [47] and Liu et al. [28] respectively propose extensions of DPO, called the multi-objective DPO (MODPO) and constrained DPO (C-DPO). A challenge common to both algorithms is the lack of flexibility regarding algorithms or datasets. Specifically, MODPO and C-DPO optimize LM policies using DPO while incorporating weighted summations of reward and safety. Hence, we must use DPO for each alignment  [^1] While it is easy to replace DPO with other alignment algorithms (e.g., KTO), MODPO or C-DPO still requires the use of the same algorithm for each metric. and prepare a dataset that contains the set of outputs $\{y\}$ characterizing both reward and safety for each prompt $\{x\}$. As an individual shortcoming, while MODPO still necessitates reward and safety modeling, C-DPO needs to iteratively apply DPO and update the Lagrangian multiplier via gradient descent.
@@ 3 Problem Formulation
Though this paper focuses on the case with a single safety function for ease of understanding, our key ideas can be easily extended to multiple safety functions. For more details, see Appendix B.
Datasets. To accommodate a wide range of situations, we relax an assumption about the dataset. We assume two different datasets exist: one for reward $\mathcal{D}_{r}$ and the other for safety $\mathcal{D}_{g}$. These two datasets do not have to share the same prompts $\{x\}$. Crucially, we do not restrict $\mathcal{D}_{r}$ and $\mathcal{D}_{g}$ to be (paired) preference datasets; that is, we accept such datasets that are used for KTO.
Example scenarios. Our problem formulation covers many real-world problems that LLMs face. Let us discuss the importance and potential usage of our formulation. One of the most direct scenarios is to reduce the harmfulness of LLMs as in Dai et al. [12]. When we define $g$ as a function to return a small value for a harmful (e.g., toxic, discriminative) answer $y$ for a given prompt $x$, our problem formulation and algorithm can be used for aligning LLMs to improve its helpfulness and harmlessness. Also, LLMs are known to be vulnerable to a variety of bias in terms of politics [26], gender [38], verbosity [36, 33], and so on. When $g$ is defined as a function to return a small value for a biased answer $y$ for a given prompt $x$, our problem formulation will help suppress such biases. A recently identified problem is that RLHF significantly reduces output diversity compared to SFT [21]. When we define $g(x,y)=-\pi(y\mid x)\log\pi(y\mid x)$, we obtain $G(\pi)=\mathbb{H}(\pi)$ for increading the diversity of an LM policy, where $\mathbb{H}$ measures policy entropy. The above are only a few examples, but our problem formulation (10) has the potential to deal with a variety of real problems LLMs face.

### v2

@@ Abstract
Safety and trustworthiness are indispensable requirements for real-world applications of AI systems using large language models (LLMs). This paper formulates human value alignment as an optimization problem of the language model policy to maximize reward under a safety constraint, and then proposes an algorithm, Stepwise Alignment for Constrained Policy Optimization (SACPO). One key idea behind SACPO, supported by theory, is that the optimal policy incorporating reward and safety can be directly obtained from a reward-aligned policy. Building on this key idea, SACPO aligns LLMs step-wise with each metric while leveraging simple yet powerful alignment algorithms such as direct preference optimization (DPO). SACPO offers several advantages, including simplicity, stability, computational efficiency, and flexibility of algorithms and datasets. Under mild assumptions, our theoretical analysis provides the upper bounds on optimality and safety constraint violation. Our experimental results show that SACPO can fine-tune Alpaca-7B better than the state-of-the-art method in terms of both helpfulness and harmlessness.
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities in diverse real-world applications [12] such as translation [52], content creation [51], coding [13, 20], summarization [41], medicine [44], and robotics [39], among others. As the utilization of LLMs in artificial intelligence (AI) systems permeates our daily lives, the importance of responsible and ethical use grows; safety issues have been highlighted [21, 29, 31]. Consequently, as AI continues to evolve and become more integrated into society, it is crucial that we actively research and develop solutions to ensure that the benefits of AI are realized while minimizing negative societal impacts.
To address this challenge, alignment [23] has been used to embed human values and goals into LLMs to enhance their utility and safety. Notably, alignment based on human feedback has emerged as a key mechanism in making LLMs more helpful and harmless, as exemplified by reinforcement learning from human feedback (RLHF, [14, 33]). Standard RLHF training flows fit a reward model to a human preference dataset and then optimize a language model (LM) policy to maximize the reward without overly diverging from the original policy. However, RLHF measures the quality of outputs in terms of a single metric (i.e., reward); thus, the achieved level of safety is not usually high, and a model that refuses to answer, while technically considered harmless, renders the response quite unhelpful [15].
@@
Safe RLHF [15] is a pioneering approach for introducing the (constrained) safe RL paradigm into the alignment of LLMs. As with the standard RLHF pipeline, Safe RLHF trains separate reward and safety models from the human preference datasets and then employs RL to optimize an LM policy. This approach facilitates the acquisition of LLMs that strike a well-balanced compromise between reward (i.e., helpfulness) and safety (i.e., harmlessness). However, the Safe RLHF pipeline is inherently more complex than standard RLHF, as it necessitates 1) fitting separate reward and safety models to preference data and then 2) learning a policy via PPO Lagrangian [36] that simultaneously optimizes an additional parameter (i.e., Lagrangian multiplier) to balance helpfulness and harmlessness. In addition, Safe RLHF often suffers from an issue called exaggerated safety behaviors [9], which results in the model generating harmless but unhelpful responses.
Our contributions. We propose an algorithm called Stepwise Alignment for Constrained Policy Optimization (SACPO) for human value alignment of LLMs while incorporating decoupled reward and safety metrics. As shown in Figure 1, SACPO is a stepwise approach that sequentially aligns an LLM with one metric (e.g., reward) and subsequently with another (e.g., safety). Our theoretical analysis allows us to employ simple RL-free alignment algorithms such as direct preference optimization (DPO, [35]) or Kahneman-Tversky optimization (KTO, [19]) for each alignment without necessitating explicit reward or safety modeling. In a theoretically justified way, SACPO allows us to use different alignment algorithms or parameters for each alignment, thereby enhancing the flexibility of the format or volume of the datasets. To enhance the practicality of SACPO, we further propose an efficient approach called P-SACPO using model merging to balance the trade-off between reward and safety performance. We provide theoretical results on the optimality and safety of the LM policy of SACPO under mild assumptions. Finally, our experimental results show that SACPO can fine-tune Alpaca-7B better than Safe RLHF in terms of both helpfulness and harmlessness.
@@ 2 Preliminaries
Given a pre-trained LLM, alignment is conventionally conducted in two stages [7, 33, 55]. In the first stage, called supervised fine-tuning (SFT), a pre-trained LLM is fine-tuned with a cross-entropy loss over high-quality human completion, resulting in a model $\pi_{\text{SFT}}$. This stage enables the model to predict the next token more properly on data more relevant for downstream tasks (e.g., dialogue, summarization). The second stage, learning from human feedback, aims to better align LLMs to human desiderata [14]. Because this paper focuses on the second stage, we will review the existing representative approaches and algorithms for the second stage. Specifically, we will briefly review RLHF [33] and then explain subsequent RL-free approaches such as DPO [35] or KTO [19].
@@ 2.1 Reinforcement Learning from Human Feedback (RLHF)
The standard RLHF pipeline consists of the following two phases: 1) reward modeling and 2) RL fine-tuning phases. With a prompt $x\in\mathcal{X}$, an LLM is regarded as a (stochastic) policy to return an output $y\in\mathcal{Y}$, where $\mathcal{X}$ and $\mathcal{Y}$ are respectively the finite spaces of prompts and outputs. Here, we assume access to a dataset of preferences $\mathcal{D}\coloneqq\{(x^{(i)},y_{w}^{(i)},y_{l}^{(i)})\}_{i=1}^{N}$, where $y_{w}$ and $y_{l}$ denote preferred and dispreferred outputs (i.e., $y_{w}\succ y_{l}$) for a prompt $x$ and $N\in\mathbb{Z}_{+}$ is the number of data. Paired outputs $(y_{w},y_{l})$ are typically sampled from $\pi_{\text{SFT}}$.
Reward modeling. In the first stage, the preference dataset $\mathcal{D}$ is assumed to be generated by a latent (unknown) reward model $r^{\star}$. A typical choice is the Bradley-Terry (BT) model [11], which stipulates that the human preference distribution $p^{\star}$ is written as
@@
RL fine-tuning. In the second stage, we aim to maximize the reward while leveraging a reverse KL divergence penalty to restrict how far the LM policy can drift from the base reference policy $\pi_{\mathrm{ref}}$, namely the SFT model $\pi_{\text{SFT}}$. Let $\pi_{\theta}$ denote the LM policy we are optimizing. We then solve the following policy optimization problem to maximize
@@
where $\rho$ is a distribution of prompts used in the RL phase, and $\mathbb{E}_{\rho,\pi}[\cdot]$ is an abbreviated notation for $\mathbb{E}_{x\sim\rho,y\sim\pi(\cdot\mid x)}[\cdot]$ for any policy $\pi\in\Pi$. Also, $\beta\geq 0$ is a hyperparameter for the KL penalty. Since this objective is not differentiable, RLHF leverages common RL algorithms such as PPO [38] as in Ouyang et al. [33] or REINFORCE [48] as in Ahmadian et al. [2] to optimize it.
@@ 2.2 Direct Learning from Human Feedback without RL
RLHF (especially when based on PPO) is computationally expensive and unstable in practice; thus, many algorithms (e.g., [35, 6, 19]) have been proposed to overcome the issues. A common idea is to analytically derive the optimal policy of (3) and parameterize it using the reward function as follows:
@@
Here, for any function $f:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ and policy $\widetilde{\pi}\in\Pi$, $Z_{f}(x;\widetilde{\pi})$ is a normalization term or constant defined as $Z_{f}(x;\widetilde{\pi})\coloneqq\sum_{y}\widetilde{\pi}(y\mid x)\exp\big{(}\frac{1}{\beta}f(x,y)\big{)}$. For the proof, see Appendix A. Note that, to derive the optimal policy (4), we do not assume a specific reward structure such as the BT model. Thus, the overall structure of the optimal policy results from the problem setting (3) characterized by the KL divergence, which is common to the representative algorithms listed below.
@@
Since $Z_{r}(x;\pi_{\mathrm{ref}})$ neither depends on $y$ nor $\pi$, by simply plugging the reward (5) into the loss function associated with the BT model (2), the resulting objective of DPO is given by
@@
As a generalized extension of DPO, Azar et al. [6] later proposed $\Psi$PO characterized by a more general objective function exclusively expressed for pairwise preferences. Also, IPO is proposed as a specific case of $\Psi$PO to make the reward function bounded to avoid overfitting.
KTO. The algorithms discussed above need a preference dataset, which is costly for humans to collect. Ethayarajh et al. [19] proposed Kahneman-Tversky optimization (KTO) that needs only a binary signal of whether the output $y$ is desirable (i.e., $y_{+}$) or undesirable (i.e., $y_{-}$) for a given prompt $x$. With an unpaired dataset $\widetilde{\mathcal{D}}\coloneqq\{(x^{(i)},y^{(i)})\}_{i=1}^{N}$, the loss function for KTO is calculated as
@@
where $v_{\text{KTO}}$ is called a value function that maps a latent reward $r_{\text{KTO}}(x,y)\coloneqq\beta\log{\frac{\pi_{\theta}(y\mid x)}{\pi_{\text{ref}}(y\mid x)}}$, relative to some reference point $\nu\coloneqq\beta\,\mathbb{D}_{\mathrm{KL}}[\,\pi_{\theta}(y\mid x)\,\|\,\pi_{\text{ref}}(y\mid x)\,]$, to its perceived value:
@@ 2.3 Safe and Multi-objective Learning from Human Feedback
Though all the algorithms discussed above consider only a singular reward function, several algorithms incorporating constraints or multiple objectives have been recently proposed [54, 15, 53, 32].
Safe RLHF. To improve the helpfulness and harmlessness of LLMs, Dai et al. [15] introduce a safety function $g^{\star}$ and then formulate the LLM alignment problem as a policy optimization task of maximizing the reward $r^{\star}$ under a safety constraint. They propose Safe RLHF that first trains reward and safety models (i.e., $r_{\psi}$ and $g_{\psi}$) using two datasets containing reward and safety information, and then solves the following problem using a popular safe RL algorithm called PPO-Lagrangian [36].
@@
Safe RLHF requires us to fit separate reward and safety models and then concurrently optimize the LM policy and Lagrangian multiplier to balance helpfulness and harmlessness. Hence, Safe RLHF is a more complex and unstable procedure, even when compared to standard RLHF.
Multi-objective and constrained DPO. Zhou et al. [54] and Liu et al. [32] respectively propose extensions of DPO, called the multi-objective DPO (MODPO) and constrained DPO (C-DPO). A challenge common to both algorithms is the lack of flexibility regarding algorithms or datasets. Specifically, MODPO and C-DPO optimize LM policies using DPO while incorporating weighted summations of reward and safety. Hence, we must use DPO for each alignment  [^1] While it is easy to replace DPO with other alignment algorithms (e.g., KTO), MODPO or C-DPO still requires the use of the same algorithm for each metric. and prepare a dataset that contains the set of outputs $\{y\}$ characterizing both reward and safety for each prompt $\{x\}$. As an individual shortcoming, while MODPO still necessitates reward and safety modeling, C-DPO needs to iteratively apply DPO while updating the Lagrangian multiplier via gradient descent.
@@ 3 Problem Formulation
Though this paper focuses on the case with a single safety function for ease of understanding, our key ideas can be easily extended to multiple safety functions. For more details, see Appendix B.1.
Datasets. To accommodate a wide range of situations, we relax an assumption about the dataset. We assume two different datasets exist: one for reward $\mathcal{D}_{r}$ and the other for safety $\mathcal{D}_{g}$. These two datasets do not have to share the same prompts $\{x\}$. Crucially, we do not restrict $\mathcal{D}_{r}$ and $\mathcal{D}_{g}$ to be (paired) preference datasets; that is, we accept such unpaired datasets that are used for KTO.
Example scenarios. Our problem formulation covers many real-world problems that LLMs face. Let us discuss the importance and potential usage of our formulation. One of the most direct scenarios is to reduce the harmfulness of LLMs as in Dai et al. [15]. When we define $g^{\star}$ as a function to return a small value for a harmful (e.g., toxic, discriminative) answer $y$ for a given prompt $x$, our problem formulation and algorithm can be used for aligning LLMs to improve their helpfulness and harmlessness. Also, LLMs are known to be vulnerable to a variety of bias in terms of politics [30], gender [42], verbosity [40, 37], and so on. When $g^{\star}$ is defined as a function to return a small value for a biased answer $y$ for a given prompt $x$, our problem formulation will help suppress such biases. A recently identified problem is that RLHF significantly reduces output diversity compared to SFT [25]. When we define $g^{\star}(x,y)=-\pi(y\mid x)\log\pi(y\mid x)$, we obtain $G(\pi)=\mathbb{H}(\pi)$ for increasing the diversity of an LM policy, where $\mathbb{H}$ measures policy entropy. The above are only a few examples, but our problem formulation (9) has the potential to deal with a variety of real problems LLMs face.